OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

he table shows the solution to the equation |2x – 3| - 1 = 2: Step 1 |2x – 3| = 2 + 1 Step 2 |2x – 3| = 3 Step 3 2x – 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct